Work out questions1.Given the following information (in millions of Birr)Government final consumption expenditure = 1400Change in stocks = 210Private final consumption expenditure = 820Net domestic capital formation = 275Exports = 175Imports = 225Depreciation = 80Factor income to abroad = 78Factor income from abroad = 194FindA.GDP by expenditure approach(2 Point)B.GNP(2 point)C.NNP(2 point)

1
Expert's answer
2022-03-13T18:53:01-0400

GNP= Final Consumption + Gross Capital Formation (investment in a firm, i.e. purchase of machinery, equipment, stocks, production sites) + government spending + net exports (exports - imports; can be either positive or negative)=1400+210+820+275+(275-175)=2805


GNP=GNP+balance of primary income exchange with the rest of the world=2805+(194-78)=2921


NNP=GNP-Depreciation=2921-80=2841